OpenStudy (anonymous):

In 1994, the average price of a new domestic car was $16,930. In 2002, the average price was $19,126. Based on a linear model, what would be the approximate predicted average price for 2008? $23,000 $21,300 $20,800 $18,600

OpenStudy (l):

2002-1994 = 8 year 2008-1994 = 16 years 16/8 = 2 19126 - 16930 = 2196 dollars in 8 years 2196 * 2 = 4392 dollars in 16 years So the price is 16930 + 4392 = 21322 dollars in 2008 *

OpenStudy (anonymous):

thank you(: your the best!!